# Om#$\mathrm{2}$

### v1

@@ 1 Introduction
We present MS-HAB [^1] Website, code, dataset, and more: https://arth-shukla.github.io/mshab, a holistic, open-sourced, home-scale manipulation benchmark with four key features: (1) fast simulation with realistic physics and manipulation, including low-level control, for efficient training, evaluation, and dataset generation, (2) home-scale manipulation tasks through the Home Assisitant Benchmark (HAB) (Szot et al., 2021), (3) extensive baselines for future work to compare against, and (4) scalable, controlled data generation using an automated, rule-based trajectory filtering system.
@@ 9 Reproducibility Statement
We open source all code for environments, training, evaluation, and data generation: https://arth-shukla.github.io/mshab. At the time of writing, the data used in this paper (490GB) is uploading to HuggingFace. While the data is uploading, users can generate data using the provided scripts in the source code. Depending on internet speed, generating may be faster than downloading.
@@ A.3 RL Subtask Evaluation Curves
A.4 Additional Experiments
A.4.1 Dataset Size
@@ A.5 Ray Tracing and Visual Fidelity
To compare performance, we run an altered version of Behavior-1k’s rendering benchmark. We use a single Nvidia RTX 3070, render 1 128x128 RGB-D image, and simulate dynamics with a simulation frequency of 120Hz and control frequency of 30Hz. Each evaluation run consists of 300 steps of random actions clipped to [-0.3, 0.3]. We report mean and 95% CIs over 10 evaluation runs.
While live-rendering with ray tracing, ManiSkill-HAB achieves $60.42\pm 0.72$ samples per second (SPS) while using $4.01\pm 0.69$ GB of GPU memory, while Behavior-1k is limited to $15.56\pm 0.04$ SPS while using $5.96\pm 0.03$ GB of GPU memory.
Hence, ManiSkill-HAB is 3.88x faster than Behavior-1k while using 32.73% less GPU memory, while also retaining similar ray-tracing render quality as seen in Fig. 10.

### v2

@@ Abstract
Figure 1: Live-rendered frames taken from ManiSkill-HAB environments while running policy rollouts with skill chaining. Ray-tracing enabled. Full videos available on website.
Videos, models, data, code, and more at http://arth-shukla.github.io/mshab
@@ 1 Introduction
We present MS-HAB [^1] Videos, models, data, code, and more: http://arth-shukla.github.io/mshab, a holistic, open-sourced, home-scale manipulation benchmark with four key features: (1) fast simulation with realistic physics and manipulation, including low-level control, for efficient training, evaluation, and dataset generation, (2) home-scale manipulation tasks through the Home Assisitant Benchmark (HAB) (Szot et al., 2021), (3) extensive baselines for future work to compare against, and (4) scalable, controlled data generation using an automated, rule-based trajectory filtering system.
@@ 9 Reproducibility Statement
We open source all code for environments, training, evaluation, and data generation, and we release our dataset for public use, all available on our website: http://arth-shukla.github.io/mshab.
@@ A.3 RL Subtask Evaluation Curves
See Appendix A.4.4 for more analysis on performance under lower collision thresholds.
@@ A.5 Ray Tracing and Visual Fidelity
To compare performance, we run an altered version of Behavior-1k’s rendering benchmark. We use a single Nvidia RTX 4090, render 1 128x128 RGB-D image, and simulate dynamics with a simulation frequency of 120Hz and control frequency of 30Hz. Each evaluation run consists of 300 steps of random actions clipped to [-0.3, 0.3]. We report mean and 95% CIs over 10 evaluation runs.
While live-rendering with ray tracing, ManiSkill-HAB achieves $69.90\pm 0.25$ samples per second (SPS) while using $6.26\pm 0.00$ GB of GPU memory, while Behavior-1k is limited to $19.92\pm 0.04$ SPS while using $7.62\pm 0.04$ GB of GPU memory.
Hence, ManiSkill-HAB is 3.51x faster than Behavior-1k while using 17.85% less GPU memory, while also retaining similar ray-tracing render quality as seen in Fig. 11.